# Common Inpainted Objects In-N-Out of Context

### v1

@@ 4.2 Verification on context classification
To verify context classification, we established strict criteria. First, an agreement was required between Molmo and annotators regarding the presence of the target object (as in Section 4.1). Second, all human annotators were to agree on the context classification (in- or out-of-context). Of our 1000 manually annotated images, a total of 477 instances met these criteria. Among these instances, Molmo’s context classification aligned with human annotations in 370 cases ($77.57\%$). These results confirm Molmo’s reliability in context understanding and, together with YOLOv8x, validate our cascaded detection and classification verification process.
5 Analysis on COinCO
With COinCO as our testbed, we explore the role of context in several tasks. We demonstrate how our data can be used to learn a general model to classify objects as in- or out-of-context. We then propose a novel Objects-from-Context task for predicting which objects naturally belong in a given scene. Finally, we evaluate state-of-the-art fake detection models and show how contextual information can enhance fake localization without any fine-tuning. These applications demonstrate the versatility of COinCO and underscore the fundamental role of context in visual understanding. Examples of our pipeline are illustrated in Figure 4.
@@ 5 Analysis on COinCO
Figure 4: Inpainting, fake detection, and objects-from-context results. Context reasoning responses are color-coded by location, size, and co-occurrence. Original objects are in red. Inpainted objects: kite, cow, horse, cup, sports ball, refrigerator, orange.
5.1 In- and out-of-context classification
COinCO’s context labels enables the training of a generalized context model to classify objects as in- or out-of-context within an image, addressing the question: “Does this object belong in this context?" We train a binary classification model that outputs a single score through a sigmoid activation function, where $1$: “out-of-context” and 00: “in-context.” For training and evaluation, we create a balanced dataset by supplementing our generated samples with original COCO images. As objects in COCO are naturally occurring, we assume they are “in-context”, ensuring a 50:50 ratio. We evaluate three models with different input modalities: (1) VisualNet: Uses only visual features from the image and mask via the VAE encoder in Stable Diffusion. (2) SemanticNet: Uses only semantic embeddings from BERT [51], including the average embedding of existing objects and the embedding of the query object. (3) VisSemanticNet: Uses both visual and semantic embeddings.
According to Table 3, the significant performance gap between VisualNet and SemanticNet shows that contextual understanding is primarily a reasoning task. VisualNet struggles to capture the complex relationships and real-world knowledge required for context reasoning and tends to predict objects as out-of-context, resulting in high recall but low accuracy and precision. VisSemanticNet achieves the best overall performance. Despite their simplicity, our context classification models show promising results and we hope they will inspire further research for more advanced context-aware models.
@@ 5.2 Objects-from-Context prediction
While our context classification model determines if an existing object belongs in a scene, our Objects-from-Context prediction task tackles a different question: “What object(s) fit to this context?” This task tests a model’s contextual comprehension of real images and aligns with the Context Challenge in [52]. We formulate this task at two levels: 1) Instance-level prediction. Predicting the exact object class among COCO’s 80 classes. 2) Clique-level prediction. Predicting the clique an object belongs to, using COCO’s super-categories: accessory, animal, appliance, electronic, food, furniture, indoor, kitchen, outdoor, person, sports, and vehicle.
We design a model that takes two inputs: the inpainted image and a binary mask indicating the target region. During training, we use the dilated bounding box of the original object as the mask and its class as the label. We use the VAE encoder from Stable Diffusion to extract latents from both inputs, which are processed by an MLP that predicts across COCO’s 80 classes. For clique-level evaluation, we map predicted object classes to their corresponding super-categories and check if they match the original objects. Importantly, our model can handle arbitrary mask sizes and locations during inference, making it adaptable for broader context reasoning applications. Detailed model architectures and training information are in the supplementary material. For comparison, we implement a baseline that ranks candidate objects based on the co-occurrence frequency of other objects in a scene [53]. If no other objects are present, this baseline defaults to random selection.
@@
Table 4 shows our model’s superior grasp of contextual features, significantly outperforming random and co-occurrence baselines. At the instance level, the substantial improvement in top-3 and top-5 accuracy (31.89% and 42.80%) indicates that even when the model’s top prediction is incorrect, the true object is often ranked among the most probable candidates. Our model demonstrates stronger performance at the clique level, correctly predicting semantically similar objects. This result demonstrates that our approach effectively learns the relationship between scene contexts and the objects that belong within them. Our model offers three key applications as shown in the three rows of Figure 5: (1) analyzing suspicious regions to predict what objects existed before manipulation, serving as a reference for anomaly detection and forensic analysis [54]; (2) suggesting contextually appropriate objects based on spatial location within a scene, enabling intelligent image editing [33]; and (3) versatile analysis of any image (including unmodified images) to identify contextually coherent objects in regions of interest, supporting context-aware content generation. This versatility advances context-driven visual understanding across multiple practical applications.
Refer to caption Refer to caption Refer to caption
Instance P(%) skis: 31.6 snowboard: 24.3 kite: 14.8 surfboard: 7.0 backpack: 5.9 Clique P(%) sports: 79.2 accessory: 6.6 person: 4.9 vehicle: 4.2 animal: 3.0 Instance P(%) kite: 22.9 skis: 22.1 snowboard: 19.6 surfboard: 8.6 person: 5.5 Clique P(%) sports: 75.1 accessory: 6.1 vehicle: 5.7 person: 5.5 animal: 4.2 Instance P(%) horse: 5.0 zebra: 5.0 giraffe: 4.9 person: 4.6 elephant: 4.4 Clique P(%) animal: 37.2 vehicle: 14.9 sports: 12.1 outdoor: 10.7 accessory: 6.8
Figure 5: Object-from-Context prediction. A red box is a query. The top row shows three examples (two inpainted, one original COCO), and the bottom row lists instance-level and clique-level predictions with their probabilities (P). Objects in red are the top predictions.
@@
5.3 Context-empowered fake localization
Fake localization aims to identify the specific synthetic regions in an image, producing a pixel-level map that is fine-grained compared to binary image-level classification. In COinCO, the ground truth mask for fake localization is the original object’s entire bounding box used in inpainting. We evaluate localization performance using four common pixel-level metrics [41, 40]: F1 Score and Accuracy (threshold of $0.5$), Area Under Curve (AUC), and Average Precision (AP).
State-of-the-art performance. We benchmark several SOTA fake detection models on COinCO (Table 5). The high accuracy and AUC scores ($>84\%$ for all models) can be attributed to the small average size of inpainted regions in our data. As most images are majority authentic, these metrics are biased towards high values. F1 and AP, which better reflect the precision of fake localization, reveal a clear performance gap among models. CAT-Net achieves the best performance across all metrics with $92.65\%$ accuracy, $76.47\%$ F1 score, $97.35\%$ AUC, and $90.33\%$ AP. Trufor obtains the second-best F1 score of $55.92\%$, while PSCC-Net shows strong performance in AUC ($95.38\%$) and AP ($79.52\%$). ManTraNet demonstrates relatively lower performance, particularly in F1 score ($30.68\%$) and AP ($50.72\%$). These findings highlight significant differences and room for improvement among SOTA fake localizers. Examples of these SOTA models’ predictions are shown in Figure 7.
Incorporating context in fake localization. To leverage contextual information for improved fake localization, we propose a simple yet effective way of enhancing the prediction scores of fake localization models in regions corresponding to out-of-context objects. Concretely, the context-enhanced prediction score $P^{\prime}(x,y)$ at pixel $(x,y)$ is defined as $\min(P(x,y)\times\gamma,\,1.0),\text{if }(x,y)\in M_{\text{OC}}$; otherwise $\text{otherwise as}P(x,y)$, the original predicted fake score, $M_{\text{OC}}$ is the mask region of out-of-context object (predicted by context reasoning), $\gamma$ is a context enhancement factor, and the $\min$ function ensures the enhanced score $\leq 1.0$.
@@ 5.3 Context-empowered fake localization
We evaluate this context-enhancement approach under two settings. In the oracle setting, we use ground truth annotations to enhance predictions within the fake object’s mask region when it is out-of-context. This serves as an upper bound for performance gains. In the practical setting, the fake objects are unknown. We propose the use of Molmo to identify suspicious objects based on size and location, deliberately excluding co-occurrence to prevent false positives. For instance, if an image contains only two objects, e.g., an apple and an inpainted traffic light, either might be seen as out-of-context. By focusing on physical violations, our method remains robust while effectively detecting fake objects. As Molmo may identify multiple out-of-context objects in an image, we define $M_{\text{OC}}$ as the union of these objects’ masks. In both settings, we set the enhancement factor $\gamma$ to $5$ to enhance prediction scores in out-of-context regions. A detailed analysis of $\gamma$ is in supplementary materials.
Figure 6 shows how our context-enhancement improves fake localization for all SOTA detectors without any finetuning. The effectiveness of our method stems from the complementary nature of context detection and fake localization. This combination ensures that even if Molmo misidentifies authentic objects as out-of-context, enhancement only strengthens the base model’s predictions, preserving robustness. These results highlight context as a valuable signal for fake detection, opening new possibilities for context-aware image forensics.
6 Conclusion and Limitations
We present COinCO, a novel dataset designed to feature in- and out-of-context objects in real scenes. By strategically replacing objects in COCO with diffusion-based inpainting, we systematically diversify the contextual status of objects in complex scenes. COinCO’s context labels allow us to train a general context classifier. We also introduce a novel Object-from-Context prediction task. Finally, our work advances current fake detection approaches by leveraging contextual information during fake pixel localization. COinCO and our findings highlight the importance of context in visual understanding and provide a foundation and testbed for future research.
While COinCO advances context-aware visual understanding, context classification remains inherently subjective despite our structured criteria, as human annotators and MLLMs occasionally disagree on contextual coherence. The Objects-from-Context prediction task is currently limited to COCO’s predefined categories. Future work could explore extending these tasks to open-vocabulary settings for more flexible context reasoning.

### v2

@@ 4.2 Verification on context classification
Following the procedure in Section 3.2, we retained 73,929 images where all three LVLMs reached unanimous agreement on the final decision, with 64,879 labeled as out-of-context and 9,050 as in-context. To validate the quality of these context annotations, we randomly sampled 1,000 images (300 in-context, 700 out-of-context) for human verification. Human annotators classified each image using the same context criteria. The LVLMs consensus achieved 95.10% agreement with human annotations, confirming the high reliability of our labels.
Among the 1,000 annotated images, we identified 164 cases where the three LVLMs agreed on the final out-of-context decision but disagreed on at least one specific criterion. Human evaluation of these disagreements revealed that Qwen2.5VL-72B achieved the highest accuracy at 87.80% (144/164). For the remaining 836 images where all LVLMs agreed on both the final decision and all criteria, human annotators assessed reasoning quality. Qwen2.5VL-72B received 94.70% preference for its detailed and interpretable explanations. Based on these validation results, we adopt Qwen2.5VL-72B’s reasoning as the primary annotations for all 73,929 images in our dataset, with the consensus from MolMo and InternVL serving to ensure annotation reliability through multi-model verification.
5 COinCO-enabled Context Tasks
With COinCO as our testbed, we explore the role of context in several tasks. We demonstrate how our data can be used to train efficient context classifiers through knowledge distillation, fine-tuning smaller models to evaluate objects based on our three context criteria. We then propose a novel Objects-from-Context task for predicting which objects naturally belong in a given scene. Finally, we evaluate state-of-the-art fake detection models and show that context can enhance fake localization without any fine-tuning. These applications demonstrate the versatility of COinCO and underscore the fundamental role of context in visual understanding. Examples of those tasks are in Figure 4.
@@ 5 COinCO-enabled Context Tasks
Figure 4: Inpainting, fake detection, and objects-from-context results. Context reasoning responses are color-coded by location, size, and co-occurrence. Original objects are in red. Inpainted objects: kite, cow, horse, cup, sports ball, refrigerator.
5.1 Fine-grained context classification
Our context classification task addresses a fundamental question: “Does an object belong in this context?” To answer this question effectively at scale, we need models that can perform accurate context reasoning while being practical for deployment. While our verification showed that Qwen2.5-VL-72B achieved excellent performance in context classification, its large size makes it impractical for many real-world deployment scenarios.
To address this challenge, we employ knowledge distillation [36, 24, 63, 53], a proven technique that transfers knowledge from large teacher models to smaller, more efficient student models. We use Qwen2.5-VL-72B as the teacher to distill its knowledge into three independent Qwen2.5-VL-3B student models—one for location, one for size, and one for co-occurrence reasoning. This distillation process reduces the model size by 24× while maintaining most of the reasoning capabilities, making context classification feasible for practical deployment. Note that this knowledge distillation task is possible thanks to the rich context reasoning labels in COinCO.
We use the 73,929 unanimous consensus images from COinCO to create balanced evaluation splits for training our distilled models. Specifically, we sample 24,000 images maintaining a 1:1 ratio of in-context to out-of-context examples, which we divide into 20,000 for training, 2,000 for validation, and 2,000 for testing. This balanced sampling ensures unbiased evaluation of context reasoning capabilities. We use context reasoning responses from QwenVL as the ground truth (discussed in Section 4.2). Importantly, each student model learns not only the final decision (in-context or out-of-context) but also the reasoning process for its specific criterion, enabling interpretable explanations.
Table 3: Context reasoning accuracy (%) on COinCO test set and original COCO images. Results on original COCO confirm models learn genuine context reasoning rather than inpainting artifacts.
COinCO Test Set Original COCO Images
Criterion Before After $\Delta$ Before After $\Delta$
Size 55.4 65.9 +10.5 52.6 83.2 +30.6
Location 67.6 75.4 +7.8 47.6 91.4 +43.8
Co-occurrence 75.5 79.9 +4.4 89.0 87.0 -2.0
Average 66.2 73.7 +7.6 63.1 87.2 +24.1
Table 3 shows consistent improvements for each student model across all criteria. On the COinCO test set, size reasoning improves most (+10.5% accuracy), while co-occurrence, despite the strongest baseline, reaches nearly 80% accuracy, confirming successful knowledge transfer from the 72B teacher to 3B students.
Validation on original COCO images. To verify that our models learn genuine semantic relationships and context rather than fitting to inpainting artifacts, we test them on 2,000 original COCO images without any manipulation. The ground-truth values of all three criteria are true as the objects are considered as in-context. The co-occurrence model achieves 89.0% accuracy on original images before fine-tuning, demonstrating it captured natural object relationships. After fine-tuning on COinCO, accuracy slightly decreases to 87.0%, suggesting the model already learned robust co-occurrence patterns that need no further adjustment. The location model shows the largest improvement after fine-tuning (+43.8% accuracy), indicating that exposure to diverse object placements in COinCO helps the model learn spatial reasoning that transfers to real images. The size model also improves substantially (+30.6% accuracy), learning appropriate scale relationships. For location and size reasoning, models achieve an average improvement of +37.2% accuracy on original COCO images, with the high performance on unmanipulated images confirming these models understand genuine contextual principles rather than memorizing dataset-specific patterns.
Each model provides interpretable natural language explanations for its decisions. Through distillation, these 3B models achieve practical inference speeds while maintaining reasoning quality [24, 62]. The modular design allows criterion-specific refinements, demonstrating that complex contextual understanding can be decomposed into focused reasoning tasks. While the 72B teacher model is too slow for practical deployment, these efficient 3B student models make context reasoning feasible for real-world applications.
Refer to caption Refer to caption Refer to caption
Instance Clique
skis 31.6 sports 79.2
snowboard 24.3 accessory 6.6
kite 14.8 person 4.9
surfboard 7.0 vehicle 4.2
backpack 5.9 animal 3.0
Instance Clique
kite 22.9 sports 75.1
skis 22.1 accessory 6.1
snowboard 19.6 vehicle 5.7
surfboard 8.6 person 5.5
person 5.5 animal 4.2
Instance Clique
horse 5.0 animal 37.2
zebra 5.0 vehicle 14.9
giraffe 4.9 sports 12.1
person 4.6 outdoor 10.7
elephant 4.4 accessory 6.8
Figure 5: Object-from-Context prediction. A red box in each image indicates the query region. The top row shows three examples (two inpainted, one original COCO). The bottom row lists instance-level and clique-level predictions with their probabilities (P(%)). Objects in red are the top predictions.
@@ 5.2 Objects-from-Context prediction
While context classification models determine if an existing object belongs in a scene, our Objects-from-Context prediction tackles a different question: “What object(s) fit this context?” This task tests a model’s contextual comprehension of real images and aligns with the Context Challenge in [55]. We formulate this task at two levels: 1) Instance-level prediction. Predicting the object class among COCO’s 80 classes. 2) Clique-level prediction. Predicting the clique an object belongs to, using COCO’s super-categories: accessory, animal, appliance, electronic, food, furniture, indoor, kitchen, outdoor, person, sports, and vehicle.
We design a model that takes two inputs: the inpainted image and a binary mask indicating the target region. During training, we use the dilated bounding box of the original object as the mask and its class as the label. We use the VAE encoder from Stable Diffusion to extract latents from both inputs, which are processed by an MLP that predicts across COCO’s 80 classes. For clique-level evaluation, we map predicted object classes to their corresponding super-categories and check if they match the original objects. Importantly, our model can handle arbitrary mask sizes and locations during inference, making it adaptable for broader context reasoning applications. Detailed model architectures and training information are in the supplementary material. For comparison, we implement a baseline that ranks candidate objects based on the co-occurrence frequency of other objects in a scene [40]. If no other objects are present, this baseline defaults to random selection. We evaluate our Objects-from-Context model on the 2,402 images inpainted from the COCO2017 validation set.
@@
Table 4 shows our model’s superior grasp of contextual features, significantly outperforming random and co-occurrence baselines. At the instance level, the substantial improvement in top-3 and top-5 accuracy (31.89% and 42.80%) indicates that even when the model’s top prediction is incorrect, the true object is often ranked among the most probable candidates. Our model demonstrates stronger performance at the clique level, correctly predicting semantically similar objects. This result demonstrates that our approach effectively learns the relationship between scene contexts and the objects that belong within them. Our model offers three key applications as shown in Figure 5: (1) analyzing suspicious regions to predict what objects existed before manipulation, serving as a reference for anomaly detection and forensic analysis [44]; (2) suggesting contextually appropriate objects based on spatial location within a scene [50, 51], enabling intelligent image editing [11]; and (3) versatile analysis of any image (including unmodified images) to identify contextually coherent objects in regions of interest, supporting context-aware content generation. This versatility advances context-driven visual understanding across multiple practical applications.
5.3 Context-empowered fake localization
Table 5: SOTA fake detection performance on COinCO.
@@ 5.3 Context-empowered fake localization
Fake localization aims to identify the specific synthetic regions in an image, producing a pixel-level map that is fine-grained compared to binary image-level classification. In COinCO, the ground truth mask for fake localization is the original object’s entire bounding box used in inpainting. We evaluate localization performance on the 2,402 test images derived from COCO2017 validation using four common pixel-level metrics [32, 39]: F1 Score and Accuracy (threshold of $0.5$), Area Under Curve (AUC), and Average Precision (AP).
State-of-the-art performance. We benchmark several SOTA fake detection models on COinCO (Table 5). The high accuracy and AUC scores ($84\%$ for all models) can be attributed to the small average size of inpainted regions in our data. As most images are majority authentic, these metrics are biased towards high values. F1 and AP, which better reflect the precision of fake localization, reveal a clear performance gap among models. CAT-Net achieves the best performance across all metrics with $92.65\%$ accuracy, $76.47\%$ F1 score, $97.35\%$ AUC, and $90.33\%$ AP. Trufor obtains the second-best F1 score of $55.92\%$, while PSCC-Net shows strong performance in AUC ($95.38\%$) and AP ($79.52\%$). ManTraNet demonstrates relatively lower performance in F1 score ($30.68\%$) and AP ($50.72\%$). These findings highlight significant differences and room for improvement among SOTA fake localizers. Examples of these SOTA models’ predictions are shown in Figure 7.
Incorporating context in fake localization. To leverage contextual information for improved fake localization, we propose a simple yet effective way of enhancing the prediction scores of fake localization models in regions corresponding to out-of-context objects. Concretely, the context-enhanced prediction score $P^{\prime}(x,y)$ at pixel $(x,y)$ is defined as $\min(P(x,y)\times\gamma,\,1.0),\text{if }(x,y)\in M_{\text{OC}}$; $\text{otherwise as}P(x,y)$, the original predicted fake score, $M_{\text{OC}}$ is the mask region of out-of-context object (predicted by context reasoning), $\gamma$ is a context enhancement factor, and the $\min$ function ensures the enhanced score $\leq 1.0$.
@@
We evaluate this context-enhancement approach under two settings. In the oracle setting, we use ground truth annotations to enhance predictions within the fake object’s mask region when it is out-of-context. This serves as an upper bound for performance gains. In the practical setting, the fake objects are unknown. We propose the use of Molmo-7B [15] to identify suspicious objects based on size and location, excluding co-occurrence to prevent false positives. For instance, if an image contains only two objects, e.g., an apple and an inpainted traffic light, either might be seen as out-of-context. By focusing on physical violations, our method remains robust while effectively detecting fake objects. As Molmo may identify multiple out-of-context objects in an image, we define $M_{\text{OC}}$ as the union of these objects’ masks. In both settings, we set the enhancement factor $\gamma$ to $5$ to enhance prediction scores in out-of-context regions. The analysis of $\gamma$ is in supplementary materials.
Figure 6 shows that our context-enhancement improves fake localization for SOTA detectors without any finetuning. The effectiveness of our method stems from the complementary nature of context detection and fake localization. Thus, even if LVLM mistakes authentic objects as out-of-context, enhancement only strengthens the base model’s predictions. Notably, out-of-context objects are not rare in real-world forgeries, especially in partial manipulation scenarios, where inserted objects often exhibit semantic inconsistencies in size, location, or scene compatibility. Our results highlight that context is valuable for fake detection.
6 Conclusion
We present COinCO, a large-scale dataset featuring both in-context and out-of-context objects created through controlled diffusion-based inpainting. By strategically replacing COCO objects, COinCO introduces diverse contextual configurations and provides rich criterion-level annotations generated through multiple LVLMs consensus. These annotations enable training efficient, interpretable context classifiers through knowledge distillation, and support a new Objects-from-Context prediction task that models what objects naturally belong in a scene. We further show that contextual cues serve as a powerful complementary signal for fake localization, improving SOTA detectors without any fine-tuning. Overall, COinCO establishes a comprehensive testbed for advancing context-aware visual understanding, contextual reasoning, and image forensics. Future work could extend context reasoning to open-vocabulary settings beyond COCO’s predefined categories, enabling more generalizable contextual understanding.
7 Acknowledgments
This work was funded in part by the National Science Foundation (SES-2521631) and the University of Georgia’s Presidential Interdisciplinary Seed Grant.